The length of a rectangle is 4 inches more than three times the width. The perimeter is 64 inches. Find the length and width.
Answer by addingup(3677)

2L + 2W = Perimeter
2(3W+4) + 2W = 64
6W+8+2W = 64
8W = 56
W = 7 this is the width
2L + 2(7) = 64
2L + 14 = 64
2L = 50
L = 25 this is your length
:
Length: 25; Width: 7